# Long and Short Guidance in Score identity Distillation for One-Step Text-to-Image Generation

### v1

@@ Abstract
Diffusion-based text-to-image generation models trained on extensive text-image pairs have shown the capacity to generate photorealistic images consistent with textual descriptions. However, a significant limitation of these models is their slow sample generation, which requires iterative refinement through the same network. In this paper, we enhance Score identity Distillation (SiD) by developing long and short classifier-free guidance (LSG) to efficiently distill pretrained Stable Diffusion models without using real training data. SiD aims to optimize a model-based explicit score matching loss, utilizing a score-identity-based approximation alongside the proposed LSG for practical computation. By training exclusively with fake images synthesized with its one-step generator, SiD equipped with LSG rapidly improves FID and CLIP scores, achieving state-of-the-art FID performance while maintaining a competitive CLIP score. Specifically, its data-free distillation of Stable Diffusion 1.5 achieves a record low FID of 8.15 on the COCO-2014 validation set, with a CLIP score of 0.304 at an LSG scale of 1.5, and a FID of 9.56 with a CLIP score of 0.313 at an LSG scale of 2. We will make our PyTorch implementation and distilled Stable Diffusion one-step generators available at https://github.com/mingyuanzhou/SiD-LSG.
@@ 2.1 Semi-implicit distributions, score identities, and model-based score matching losses
It adopts $f_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ as the functional approximation of the conditional expectation of the real image $\boldsymbol{x}_{0}$ given noisy image $\boldsymbol{x}_{t}$ and text ${\boldsymbol{c}}$, expressed as $\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t},{\boldsymbol{c}}]$, adopts $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},{\boldsymbol{c}})$ to predict the noise inside $\boldsymbol{x}_{t}$, and adopts $-\sigma_{t}^{-1}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},{\boldsymbol{c}})$ as the functional approximation of the true score $\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}})$. Given time step $t\sim p(t)$ and text ${\boldsymbol{c}}$, we define the model-based explicit score-matching (MESM) distillation loss, which is a form of Fisher divergence (Lyu, 2009; Holmes and Walker, 2017; Yang et al., 2019; Yu and Zhang, 2023), as
@@ 2.2 SiD with classifier-free guidance
    Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and let $\boldsymbol{x}_{g}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$; Sample $t\in\{t_{\min},t_{\min}+1,\ldots,t_{\max}\}$, compute $\omega_{t}$ with (9), and let $\boldsymbol{x}_{t}=\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$
@@
$=\textstyle\frac{\omega(t)}{{\sigma_{t}^{2}}}(\boldsymbol{\epsilon}_{\psi,\kappa_{4}}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})-\boldsymbol{\epsilon}_{\phi,\kappa_{2}}(\boldsymbol{x}_{t},t,{\boldsymbol{c}}))^{T}(\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\psi,\kappa_{3}}(\boldsymbol{x}_{t},t,{\boldsymbol{c}}))$
@@ Appendix E Prompts and additional example images
    chinese red blouse, in the style of dreamy and romantic compositions, floral explosions.
@@
    poster art for the collection of the asian woman, in the style of gloomy, dark orange and white, dynamic anime, realistic watercolors, nintencore, weathercore, mysterious realism.
@@
    very beautiful girl in bright leggings, white short top, charismatic personality, professional photo, style of jessica drossin, super realistic photo, hyper detail, great attention to skin and eyes, professional photo.
@@
    pumpkins, autumn sunset in the old village, cobblestone houses, streets, plants, flowers, entrance, realistic, stunningly beautiful.
@@
    Highly detailed mysterious egyptian (sphynx cat), skindentation:1.2, bright eyes, ancient egypt pyramid background, photorealistic, (hyper-realistic:1.2), cinematic, masterpiece:1.1, cinematic lighting.
@@
    vw bus, canvas art, abstract art printing, in the style of brian mashburn, light red and light brown, theo prins, charming character illustrations, pierre pellegrini, vintage cut-and-paste, rusty debris.
@@
    painterly style, seductive female League of legends Jinx character fighting at war, raging, crazy smile, crazy eyes, rocket lancher, guns, crazy face expression, character design, body is adorned with glowing golden runes, intense green aura around her, body dynamic epic action pose, intricate, highly detailed, epic and dynamic composition, dynamic angle, intricate details, multicolor explosion, blur effect, sharp focus, uhd, hdr, colorful shot, stormy weather, tons of flying debris around her, dark city background.
@@
    A watercolor painting of a vibrant flower field in spring, with a rainbow of blossoms under a bright blue sky. 8K, best quality, fine details.
@@
    digital art of a beautiful tiger pokemon under an apple tree, cartoon style,Matte Painting,Magic Realism,Bright colors,hyper quality,high detail,high resolution.
@@
    painterly style, Goku fighting at war, raging, blue hair, character design, body is adorned with glowing golden runes, yellow aura around him, body dynamic epic action pose, intricate, highly detailed, epic and dynamic composition, dynamic angle, intricate details, multicolor explosion, blur effect, sharp focus, uhd, hdr, colorful shot, stormy weather, tons of flying debris around him, dark city background.
@@
    A stunning steampunk city with towering skyscrapers and intricate clockwork mechanisms, gears and pistons move in a complex symphony, steam billows from chimneys, airships navigate the bustling skylanes, a vibrant metropolis.
@@
    Samurai looks at the enemy, stands after the battle, fear and horror on his face, tired and beaten, sand on his face mixed with sweat, an atmosphere of darkness and horror, hyper realistic photo, In post - production, enhance the details, sharpness, and contrast to achieve the hyper - realistic effect.
@@
    A portrait of an elemental entity with strong rim lighting and intricate details, painted digitally by Alvaro Castagnet, Peter Mohrbacher, and Dan Mumford.
@@
    A regal female portrait with an ornate headdress decorated with colorful gemstones and feathers, her robes rich with intricate designs and bright hues. 8K, best quality, fine details.
@@
    A detailed painting of Atlantis by multiple artists, featuring intricate detailing and vibrant colors.
@@
    A landscape featuring mountains, a valley, sunset light, wildlife and a gorilla, reminiscent of Bob Ross’s artwork.

### v2

@@ Abstract
Diffusion-based text-to-image generation models trained on extensive text-image pairs have shown the capacity to generate photorealistic images consistent with textual descriptions. However, a significant limitation of these models is their slow sample generation, which requires iterative refinement through the same network. In this paper, we enhance Score identity Distillation (SiD) by developing long and short classifier-free guidance (LSG) to efficiently distill pretrained Stable Diffusion models without using real training data. SiD aims to optimize a model-based explicit score matching loss, utilizing a score-identity-based approximation alongside the proposed LSG for practical computation. By training exclusively with fake images synthesized with its one-step generator, SiD equipped with LSG rapidly improves FID and CLIP scores, achieving state-of-the-art FID performance while maintaining a competitive CLIP score. Specifically, its data-free distillation of Stable Diffusion 1.5 achieves a record low FID of 8.15 on the COCO-2014 validation set, with a CLIP score of 0.304 at an LSG scale of 1.5, and a FID of 9.56 with a CLIP score of 0.313 at an LSG scale of 2. Our code and distilled one-step text-to-image generators are available at https://github.com/mingyuanzhou/SiD-LSG.
distinguished older gentleman in a vintage study, surrounded by books and dim lighting […]
saharian landscape at sunset , 4k ultra realism, BY Anton Gorlin, trending on artstation, […]
chinese red blouse, in the style of dreamy and romantic compositions, floral explosions. […]
futuristic simple multilayered architecture, habitation cabin in the trees, dramatic soft light […]
poster art for the collection of the asian woman, dynamic anime […], mysterious realism, […]
A fantasy-themed portrait of a female elf with golden hair and violet eyes, her attire […]
very beautiful girl in […], white short top, charismatic personality, professional photo, […]
steampunk atmosphere, a stunning girl with a mecha musume aesthetic, adorned in […]
tshirt design, colourful, no background, yoda with sun glasses, dancing at a festival […] 8k.
@@ 2.1 Semi-implicit distributions, score identities, and model-based score matching losses
It adopts $f_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ as the functional approximation of the conditional expectation of the real image $\boldsymbol{x}_{0}$ given noisy image $\boldsymbol{x}_{t}$ and text ${\boldsymbol{c}}$, expressed as $\mathbb{E}[\boldsymbol{x}_{0}\,|\,\boldsymbol{x}_{t},{\boldsymbol{c}}]$, adopts $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})$ to predict the noise inside $\boldsymbol{x}_{t}$, and adopts $-\sigma_{t}^{-1}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},{\boldsymbol{c}})$ as the functional approximation of the true score $\nabla_{\boldsymbol{x}_{t}}\ln p_{\text{data}}(\boldsymbol{x}_{t}\,|\,{\boldsymbol{c}})$. Given time step $t\sim p(t)$ and text ${\boldsymbol{c}}$, we define the model-based explicit score-matching (MESM) distillation loss, which is a form of Fisher divergence (Lyu, 2009; Holmes and Walker, 2017; Yang et al., 2019; Yu and Zhang, 2023), as
@@ 2.2 SiD with classifier-free guidance
    Sample $\boldsymbol{z}\sim\mathcal{N}(0,\mathbf{I})$ and ${\boldsymbol{c}}$, and let $\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z},{\boldsymbol{c}})=f_{\theta}(\sigma_{t_{\text{init}}}\boldsymbol{z},t_{\text{init}},{\boldsymbol{c}})$; Sample $t\in\{t_{min},\ldots,t_{\max}\}$ and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$, compute $\omega_{t}$ with (9), and let $\boldsymbol{x}_{t}=a_{t}\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t}$
@@
$=\textstyle\frac{\omega(t)}{{\sigma_{t}^{2}}}(\boldsymbol{\epsilon}_{\psi,\kappa_{2}}(\boldsymbol{x}_{t},t,{\boldsymbol{c}})-\boldsymbol{\epsilon}_{\phi,\kappa_{4}}(\boldsymbol{x}_{t},t,{\boldsymbol{c}}))^{T}(\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\psi,\kappa_{3}}(\boldsymbol{x}_{t},t,{\boldsymbol{c}}))$
@@ Appendix E Prompts and additional example images
    chinese red blouse, in the style of dreamy and romantic compositions, floral explosions –ar 24:37 –stylize 750 –v 6
@@
    poster art for the collection of the asian woman, in the style of gloomy, dark orange and white, dynamic anime, realistic watercolors, nintencore, weathercore, mysterious realism –ar 69:128 –s 750 –niji 5
@@
    ’very beautiful girl in bright leggings, white short top, charismatic personality, professional photo, style of jessica drossin, super realistic photo, hyper detail, great attention to skin and eyes, professional photo.
@@
    pumpkins, autumn sunset in the old village, cobblestone houses, streets, plants, flowers, entrance, realistic, stunningly beautiful
@@
    "Highly detailed mysterious egyptian (sphynx cat), skindentation:1.2, bright eyes, ancient egypt pyramid background, photorealistic, (hyper-realistic:1.2), cinematic, masterpiece:1.1, cinematic lighting"
@@
    "vw bus, canvas art, abstract art printing, in the style of brian mashburn, light red and light brown, theo prins, charming character illustrations, pierre pellegrini, vintage cut-and-paste, rusty debris –ar 73:92 –stylize 750 –v 6"
@@
    painterly style, seductive female League of legends Jinx character fighting at war, raging, crazy smile, crazy eyes, rocket lancher, guns, crazy face expression, character design, body is adorned with glowing golden runes, intense green aura around her, body dynamic epic action pose, intricate, highly detailed, epic and dynamic composition, dynamic angle, intricate details, multicolor explosion, blur effect, sharp focus, uhd, hdr, colorful shot, stormy weather, tons of flying debris around her, dark city background, modifier=CarnageStyle, color=blood_red, intensity=1.6
@@
    "A watercolor painting of a vibrant flower field in spring, with a rainbow of blossoms under a bright blue sky. 8K, best quality, fine details.",
@@
    "digital art of a beautiful tiger pokemon under an apple tree, cartoon style,Matte Painting,Magic Realism,Bright colors,hyper quality,high detail,high resolution, –video –s 750 –v 6.0 –ar 1:2"
@@
    "painterly style, Goku fighting at war, raging, blue hair, character design, body is adorned with glowing golden runes, yellow aura around him, body dynamic epic action pose, intricate, highly detailed, epic and dynamic composition, dynamic angle, intricate details, multicolor explosion, blur effect, sharp focus, uhd, hdr, colorful shot, stormy weather, tons of flying debris around him, dark city background, modifier=CarnageStyle, color=blood_red, intensity=1.6"
@@
    A stunning steampunk city with towering skyscrapers and intricate clockwork mechanisms, gears and pistons move in a complex symphony, steam billows from chimneys, airships navigate the bustling skylanes, a vibrant metropolis
@@
    "Samurai looks at the enemy, stands after the battle, fear and horror on his face, tired and beaten, sand on his face mixed with sweat, an atmosphere of darkness and horror, hyper realistic photo, In post - production, enhance the details, sharpness, and contrast to achieve the hyper - realistic effect"
@@
    A portrait of an elemental entity with strong rim lighting and intricate details, painted digitally by Alvaro Castagnet, Peter Mohrbacher, and Dan Mumford
@@
    "A regal female portrait with an ornate headdress decorated with colorful gemstones and feathers, her robes rich with intricate designs and bright hues. 8K, best quality, fine details.",
@@
    "A detailed painting of Atlantis by multiple artists, featuring intricate detailing and vibrant colors.",
@@
    "A landscape featuring mountains, a valley, sunset light, wildlife and a gorilla, reminiscent of Bob Ross’s artwork.
A fantasy-themed portrait of a female elf with golden hair and violet eyes, […]
pumpkins, autumn sunset in the old village, cobblestone houses, streets, […]
Highly detailed mysterious egyptian (sphynx cat), skindentation:1.2, […]
vw bus, canvas art, abstract art printing, in the style of brian mashburn […]
painterly style, seductive female League of legends Jinx character fighting at […]
A charismatic chef in a bustling kitchen, his apron dusted with flour, smiling […]
A young adventurer with tousled hair and bright eyes, wearing a leather […]
A watercolor painting of a vibrant flower field in spring, with a rainbow of […]
digital art of a beautiful tiger pokemon under an apple tree, cartoon style, […]
painterly style, Goku fighting at war, raging, blue hair, character design, body […]
A stunning steampunk city with towering skyscrapers and intricate […]
Samurai looks at the enemy, stands after the battle, fear and horror on his […]
A portrait of an elemental entity with strong rim lighting and intricate […]
A regal female portrait with an ornate headdress decorated with gemstones[…]
A detailed painting of Atlantis by multiple artists, featuring intricate […]
A landscape featuring mountains, a valley, sunset light, wildlife and a […]